find the discriminant , and determine the number of real solutions. Then solve 3x^2+2x-6=0
Answer by AnlytcPhil(1276)

find the discriminant , and determine the number of real 
solutions. Then solve 3x^2+2x-6=0

a=3, b = 2, c = -6

b²-4ac = 2²-4(3)(-6) = 4+72 = 76

The discriminant is 76, which is a positive number so
there are 2 real solutions.

If it had been 0, there would have been 1 real solution.

It it had been a negative number, there would have been 0 
(no) real solution.

Edwin